OpenStudy (anonymous):

The area of a circle is 50 find the radius

OpenStudy (anonymous):

Area of a circle=pi times the radius squared \[A=\pi*r^2\]

OpenStudy (anonymous):

\[50=\pi*r^2\]

OpenStudy (anonymous):

\[50/\pi=r^2\]

OpenStudy (anonymous):

\[\sqrt{50/\pi}=r\]

OpenStudy (anonymous):

I got 2.13 but it isnt right :/

OpenStudy (anonymous):

I got 3.989

OpenStudy (anonymous):

just divide 50 by pi and then take the square root of that answer.

OpenStudy (anonymous):

what does 15.92356688=?

OpenStudy (anonymous):

that is 50/pi

OpenStudy (anonymous):

you still need to take the square root